OpenStudy (anonymous):

help?!?!(medal will be given) Use the following four coordinates to determine which one has a distance of 5 units from point A (2, 6). (5, 2) (2, 5) (6, 7) (7, 5)

OpenStudy (anonymous):

i think you are going to get lucky and get it on the first try

OpenStudy (anonymous):

\[(2, 6), (5, 2) \] from \(2\) to \(5\) is \(5-2=3\) units to the right from \(6\) to \(2\) is \(6-2=4\) units down by pythagoras the square of the distance is \[3^2+4^2=9+16=25\] making the distance \(\sqrt{25}=5\)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw